4. Find the domain of f(x) V19x-38 Give your answer in interval notation
Answer #1

Domain is set of all real value of x which makes the function to work and will generate real output f(x)

Given

F(x) =  ?(19x -38)

So domain of F(x) is

All real value of x : x \geq 2

In interval form

Domain = \left (-\infty \left. \lef ,2 \right ]